OpenStudy (anonymous):

find all x-coordinates of all the relative extreme & inflection points of f(x) = 1/3Px^3 + 2sqrt(P)x^2 - 5x +2. Where p is greater that 0 & is a constant. We were told not to substitut a number in for P. I just am not sure if i am doing it right, I found first derivative of = px^2 + 4sqrt(p)x - 5 = and having a very hard time knowing if i did this right = Critical Points = (-4sqrt(p))/2p Also for inflection points I took the second derivative and found it to = 2px+ 4sqrt(p) & the zero of that I found to be = -2/sqrt(p) = inflection points - am I correct at all on these or way off?

OpenStudy (aum):

f(x) = 1/3Px^3 = 2sqrt(P)x^2 - 5x +2 Is the second equal sign supposed to be + ?

OpenStudy (anonymous):

Yes, so sorry

OpenStudy (anonymous):

Just fixed that

OpenStudy (aum):

\[ f(x) = \frac 13Px^3 + 2\sqrt{P}x^2 - 5x +2 \\ f'(x) = Px^2 + 4\sqrt{P}x - 5 = 0 \\ x = \frac{-4\sqrt{P} \pm \sqrt{16P + 20P}}{2P} = \frac{-4\sqrt{P} \pm 6\sqrt{P}}{2P} = \frac{-5\sqrt{P}}{P}, ~~~\frac{\sqrt{P}}{P}\\ \]

OpenStudy (anonymous):

Thank you! I was doing the quad equation and forgot to square the stinkin 4 as well and was coming up with weird answers. Is my critical point correct?

OpenStudy (aum):

The two x values shown above are the critical points:\[ x = \frac{-5\sqrt{P}}{P}, ~~~x = \frac{\sqrt{P}}{P}\\ \]

OpenStudy (anonymous):

Sorry, I meant inflection points

OpenStudy (aum):

\[ f''(x) = 2Px + 4\sqrt{P} = 0 \\ x = -\frac{-4\sqrt{P}}{2P} = -2\frac{\sqrt{P}}{P} \]

OpenStudy (anonymous):

Can you simplify that to -2/sqrt(p)? or not?

OpenStudy (aum):

Many textbooks and teachers don't like a radical in the denominator. They ask you to rationalize the denominator.

OpenStudy (anonymous):

Ok, but it is the same thing? Just making sure i understand that math, and my teacher is a little bizzar

OpenStudy (aum):

They are identical.

OpenStudy (anonymous):

Thanks so much for your help!!

OpenStudy (aum):

They are identical. But some teachers and online test software mark the answer wrong if the denominator is not rationalized. So I will play it safe expressing the answers for both the critical points and the inflection point in the form I did with no radicals in the denominator. You are welcome.